OpenStudy (anonymous):

Please help. Last chance to get my diploma In a right triangle, the acute angles measure x + 15 and 2x degrees. What is the measure of the larger acute angle of the triangle? 15° 25° 40° 50°

OpenStudy (anonymous):

I think its 50 but don't count on me.

OpenStudy (anonymous):

How did you get it?

OpenStudy (anonymous):

I think i got it wrong. hold on.

OpenStudy (anonymous):

Nevermind, its 50. The angles in a triangle sum to 180 degrees. So as the right angle is 90, the other two must total 180 - 90 = 90. So, x+15 + 2x = 90 Therefore 3x = 90 - 15 = 75 and so x=25. So the smaller angles are x+15 = 40 degrees and 2x=50 degrees. So the larger one is 50 degrees.

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

Your welcome